OpenStudy (ksaimouli):

15pie=r*(15pie/6) plz slove

OpenStudy (karatechopper):

want help?

OpenStudy (ksaimouli):



OpenStudy (anonymous):

multiply both sides by 6/15pi 6=r

OpenStudy (ksaimouli):

yup

OpenStudy (karatechopper):

lets take this piece by piece

OpenStudy (ksaimouli):

but the answer is 18

OpenStudy (karatechopper):

hold on there!!

OpenStudy (karatechopper):

step by step..

OpenStudy (karatechopper):

15 x 3.14 is.....

myininaya (myininaya):

\[15 \pi =r \frac{15 \pi }{6} => \frac{15 \pi \cdot 6}{15 \pi} =r => r=6\]